Anle Corporation has a current price of , is expected to pay a dividend of

in one year, and its expected price right after paying that dividend is .

a. What is Anle's expected dividend yield?

b. What is Anle's expected capital gain rate?

c. What is Anle's equity cost of capital?
Answer #1

A) 2/14= 0.1428571429 or answer: 14.29%


B) (24-14)/14= 0.7142857143 or answer: 71.43%


C) 14.29+ 71.43= answer: 85.72

answered by: Andrew San Andres